# On the attribution of weather events to climate change using empirically-fit extreme value distributions

### v3

@@ Abstract
Changes in extreme weather events are a potentially important aspect of anthropogenic climate change (ACC), yet, are difficult to attribute to ACC because the record length is often similar to, or shorter than, extreme-event return periods. This study is motivated by the “World Weather Attribution Project” (WWA) and, specifically, their approach of fitting extreme value distribution functions to local observations. The approach calculates the dependence of distribution parameters on global mean surface temperature (GMST) and uses this dependence to attribute extreme events to ACC. Applying this method to preindustrial climate simulations with no time-varying greenhouse gas forcing, we still find a strong dependence of distribution parameters on GMST. This dependence results from internal climate variability (e.g., ENSO) affecting both extreme events and GMST. Therefore, dependence on GMST does not necessarily imply an effect of ACC on extremes. We further consider for these three cases whether an extreme value, normal, or log-normal distribution better represents the data; if a GMST-dependence of distribution parameters is justified using a likelihood ratio test; and if a meaningful attribution is possible given uncertainties in GMST dependence. In one case we find that an attribution of Australia’s 2020–2021 Bushfires to ACC is difficult due to the effects of internal variability. In another involving the 2019–2021 drought in Madagascar we find that the small number of available data points precludes a meaningful attribution analysis. Overall, we find that the effects of internal climate variability on GMST and the uncertain relationship between GMST and regional extremes may lead to inaccurate attribution conclusions using the WWA approach.
@@ 1 Introduction
One particular approach that is the focus of this paper and that has been used in recent years by the WWA as a rapid response to extreme events attempts to fit the observed data to an extreme value PDF The World Weather Attribution project (2024). The fit allows for a dependence of the distribution parameters on explanatory variables, typically the GMST (e.g., Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022; Philip et al., 2020). A GMST dependence is then interpreted as a response of extreme events to ACC. This dependence is then used to attribute extreme events by quantifying the effects of anthropogenic forcing on the probability of their occurrence. Technically, this is achieved by writing the location parameter of the extreme event distribution, $\mu$, as a function of $\overline{T}$, the (low-pass filtered, via a 4-year running average) GMST “as a measure of anthropogenic climate change” Harrington et al. (2022), such that $\mu=\mu_{0}+\alpha\overline{T}$. The GMST-dependence, $\alpha$, is estimated via a fit to observations. The fitted PDF is then used to evaluate the probability of extreme events at 1900, $P(\mu_{0}+\alpha\overline{T}(1900))$ relative to those at 2020, $P(\mu_{0}+\alpha\overline{T}(2020))$. This allows for estimating how the probability of an extreme event changed between these times. A change in the probability of an extreme event is inferred by the WWA approach if $\alpha$ is non-zero and in proportion to the change in GMST.
Our focus is on the empirically-based approaches used by the WWA, but it should be noted that WWA also requires evidence from climate model simulations before making any attribution statements. WWA often uses multiple climate models and is careful to exclude models based on their performance and return periods relative to observations (e.g., Pinto et al., 2023). In this study we use a single model (CESM) to demonstrate some of our points but focus on assessment of observational methodologies associated with fitting of extreme value distributions to observations. Our goal is limited to evaluating the observational approach described by WWA and suggesting additional tests to examine its robustness, as opposed to developing alternative attribution approaches.
@@
The issue exposed by our analysis is simple: internal variability modes such as ENSO affect both local temperature extremes and GMST. In some cases, the effect on local extremes (say the effects of ENSO on heatwaves in Australia) may be large, but the effect on GMST is small. This might lead to a false conclusion that a small GMST signal leads to a comparatively large effect on local extremes. To demonstrate this point, we analyze a preindustrial climate model run with no anthropogenic forcing. We find, for example, that the fit of Australian heat extremes leads to values of $\alpha$ that are substantially greater than zero in the absence of ACC because of the effects of internal climate variability modes such as ENSO on both GMST and on local extremes. Our findings show that $\alpha$ can be strongly affected by internal variability, biasing the attribution results and complicating the WWA approach.
@@ 2.1 The Generalized Extreme Value (GEV) distribution
Extreme value distributions are a family of PDFs that can be used to represent the statistical behavior of block maxima or minima of a record. An example is an annual time series of the maximum daily temperature at each year, the “block” being one year in this case. It can be shown that the distribution of such maxima follows one of three classes of PDFs: Gumbel, Fréchet, or Weibull Coles et al. (2001), depending on the shape of the tail of the distribution of the events for which block maxima are calculated. Rather than explicitly specifying one of these three classes in fitting extreme data, one can combine these distributions into one functional form, represented by the Generalized Extreme Value (GEV) distribution. Large quantities of data are typically needed to accurately fit a GEV and calculate its parameters (e.g., Philip et al., 2020; Trevino et al., 2020).
@@ 2.2 Generalized Pareto Distribution (GPD)
The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/u_{0})$ applied to both $u_{0}$ and $\sigma_{0}$ is strictly positive and maintain a fixed ratio. We obtain the optimized GPD parameters ($u_{0},\sigma,\xi,\alpha$) by again maximizing the log-likelihood (Eq. 4.10, Coles et al., 2001),
@@ 2.3 Uncertainty estimates and empirical cumulative distribution functions
We follow WWA studies and use non-parametric bootstrapping to estimate uncertainty ranges for the optimized parameters. In particular, we use the uncertainty in $\alpha$ as one way to evaluate the validity of the attribution results. We also compare the appropriately fitted distributions (GEV and normal distributions for extreme temperatures or GPD and log-normal for droughts) to the empirical cumulative distribution functions (CDFs) of the observations to assess the quality of the fit. The empirical CDF is calculated by sorting extreme values (either temperature or precipitation) from smallest to largest, calculating the cumulative sum, and normalizing to a maximum CDF value of one. We calculate error bars in the empirical CDFs based on bootstrap resampling with replacement of the observations. The empirical CDF is calculated for each sampling, and we then calculate the 90th and 95th percentiles from the resulting distribution at each value of the variable whose CDF is estimated. In order to estimate uncertainty in the parameters estimated from our GEV or GPD maximum likelihood fit, we similarly use bootstrap resampling. We sample with replacement the GMST and corresponding extreme temperature record 5,000 times and optimize the parameters for each resampling. We then estimate the 90th and 95th percentiles from 5,000 bootstrap samples. For the Madagascar precipitation case, we perform 5,000 bootstrap resamples when analyzing the observations, but 1,000 bootstrap resamples when analyzing the CESM Large Ensemble due to the slow convergence and expensive computations using the GPD log-likelihood maximization.
@@ 2.6 Hypothesis testing
Although WWA does not provide a detailed discussion of the hypothesis that they assume, we interpret that they adopt a null hypothesis of $\alpha$ equal to zero implying no ACC influence on extremes and an alternative hypothesis that, if $\alpha$ is not equal to zero, ACC does influence extremes. This inference is based on the quote, “If the one- or two-sided 95% confidence interval [for the change in probability, which has the same effect as $\alpha$] excludes zero, the change is statistically significant” (see Section 4.3.1 of Philip et al., 2020). We assess whether rejection of a null hypothesis of $\alpha$ equal to zero permits accepting an alternate of ACC influencing extreme events. In other words, we assess if a nonzero $\alpha$ necessarily means a role for ACC. We note that another null hypothesis for purposes of heat could be that the location of PDF of heat extremes is shifted by an amount equaling the GMST warming, as would be represented by $\alpha=1$. For purposes of consistency with the null implied by WWA, however, we only consider a null of $\alpha$ equal to zero.
WWA uses two variants of their analysis approach, one in which the extreme data in question is incorporated into the estimate of the PDFs, which we follow here, and one in which it is excluded Philip et al. (2020). The act of choosing to study a class of events immediately following the occurrence of a particularly extreme example is known to induce a bias related to whether it is included or excluded (Miralles and Davison, 2023, see). We choose to include the extreme value in question, which tends to make the parametric dependence on GMST larger and tends to make it easier to reject the null of no change in favor of the alternate hypothesis. Within the context of the WWA hypothesis test, this choice to include the extreme in question makes it more likely to conclude that AAC significantly influences extreme events.
@@ 2.7 Data
We also study GMST anomalies and extreme temperatures over the Siberian region defined in CC2021 within the CESM Large Ensemble Kay et al. (2015) over the period 1926–2019. The CESM Large Ensemble contains 40 simulations of the climate over the period 1920–2100 under historical and RCP 8.5 external forcing, of which we use the 35 that were run on the NCAR supercomputer and not those runs done on the University of Toronto supercomputer that contain a global mean temperature bias National Center for Atmospheric Research (2024). These simulations differ only in small perturbations to the initial conditions, which lead to different temporal sequences of modes of internal variability. We use model output only up to the year 2020, and, as a result, there is little sensitivity to the emissions scenario used (DAMIP, Gillett et al., 2016, uses RCP 4.5) because emissions scenarios only significantly diverge later.
@@ 3 Results
We first describe our analyses of how extreme events depend on GMST in the context of the long CESM pre-industrial simulation in subsection 33.1. The simulation contains no time-varying anthropogenic forcing, making it useful for evaluating a null-hypotheses associated with no anthropogenic influence. We then consider each of the three case studies in subsections 33.2, 33.3, and 33.4. For each case we assess how GEV or GPD distributions fit the observations relative to normal or log-normal distributions. We then evaluate evidence for the inclusion of a GMST dependence within distributions using the deviance statistic and estimate the uncertainty level of the parameter $\alpha$ used for the attribution. Finally, we examine the amount of data needed to adequately constrain GMST dependence in a GEV or GPD fit.
@@ 3.1 Extremes analysis of a preindustrial simulation
WWA recognizes that $\alpha$ may be influenced by internal variability and, as noted, use a 4-year running mean GMST to filter out the effects of short-term internal variability. This approach does not fully remove the ENSO signal, however, because much of the variability is at longer periods. Spectral analysis of the annual-average Bivariate ENSO time series Smith and Sardeshmukh (2000) between 1870–2022, for example, indicates that 63% of the variance in the time series resides at frequencies below 1/(4 years). There is also the potential for nonlinear interactions between ENSO and GMST and the presence of slower modes of variability such as the Pacific Decadal Oscillation (PDO).
@@
A similar analysis of Australian heat extremes made using the CESM Large Ensemble, which includes anthropogenic forcing, gives a distribution of $\alpha$ that is closer to zero (blue bars in Fig. 1) than in the unforced simulations (green bars). This indicates that both ACC and internal variability influence regional extremes and GMST, such that estimates of $\alpha$ will reflect multiple processes. In simulations with time-varying anthropogenic forcing, however, there is no way to disentangle the effect of internal variability on $\alpha$ using the WWA approach as described in published literature. We conclude that the magnitude of $\alpha$ is a biased indicator of anthropogenic influence. Including indices of internal variability as covariates may help to disentangle their respective contributions, though the projection of internal variability onto GMST can still confound the results.
@@
These results highlight a potential major difficulty with analyses based on a fit of distribution parameters to GMST in a more realistic scenario that includes both internal variability and ACC: the calculated $\alpha$, and therefore the shift in the probability of extreme events, may be due to internal variability rather than only ACC. When analyzing the observational record in section 3.3, we find $\alpha\approx 2\;^{\circ}$ per C. This suggests a two-degree warming of extreme events for a one-degree warming of the GMST. Based on our results from the CESM preindustrial run with a time-invariant anthropogenic forcing, however, we suspect that the estimate obtained for $\alpha$ from the observed record is biased high by internal variability.
@@
Analysis of the CESM preindustrial run for the case of Siberian heatwaves (green bars in Fig. 1b) shows that the value for $\alpha$ deduced from a record similar in length to that of the observations ranges between 1 to 4, although there is again no time-varying anthropogenic forcing in the corresponding model run. This large and variable response implies that a non-zero $\alpha$ does not necessarily imply an effect of ACC. This analysis shows that $\alpha$ equal to zero is an unsuitable null hypothesis for the purposes of evaluating the presence of an anthropogenic influence on extreme events.
@@
The result that internal variability may strongly affect $\alpha$ when the 4-year average GMST is used to link extreme events to ACC suggests that it might be useful to reevaluate some of the WWA results based on the fit to extreme distributions in at least the three studies that we review. It is possible, of course, that a strong anthropogenic signal in a few decades could significantly affect extreme events and overwhelm the effects of internal variability in cases such as the Siberian Heatwave. The attribution of extreme temperatures may be more difficult in cases like the Australian example, where the effects of internal variability are large, as indicated by the large $\alpha$ detected in the CESM preindustrial simulation. In such cases, it would be important to correctly identify all relevant covariates — such as the possible effect of ENSO in the Australian case — and note that misleading results for $\alpha$ may be obtained if one relies only on GMST.

### v4

@@ Abstract
Changes in extreme weather events are a potentially important aspect of anthropogenic climate change (ACC), yet, are difficult to attribute to ACC because the record length is often similar to, or shorter than, extreme-event return periods. This study is motivated by the “World Weather Attribution” initiative (WWA) and, specifically, their approach of fitting extreme value distribution functions to local observations. They calculate the dependence of distribution parameters on global mean surface temperature (GMST) and use this dependence to attribute extreme events to ACC. Applying this method to preindustrial climate simulations with no time-varying greenhouse gas forcing, we still find a strong dependence of distribution parameters on GMST. This dependence results from internal climate variability (e.g., ENSO) affecting both extreme events and GMST. Therefore, dependence on GMST does not necessarily imply an effect of ACC on extremes. We further consider whether an extreme value, normal, or log-normal distribution better represents the data; if a GMST-dependence of distribution parameters is justified using a likelihood ratio test; and if a meaningful attribution is possible given uncertainties in GMST dependence. We find, for example, that an attribution of Australia’s 2020–2021 Bushfires to ACC is difficult due to the effects of internal variability. For the 2019–2021 drought in Madagascar we find that the small number of available data points precludes a meaningful attribution analysis. Overall, we find that the effects of internal climate variability on GMST and the uncertain relationship between GMST and regional extremes may lead to inaccurate attribution conclusions using the part of the WWA approach examined here.
@@ 1 Introduction
One particular approach that is the focus of this paper and that has been used in recent years by WWA as a rapid response to extreme events attempts to fit the observed data to an extreme value PDF The World Weather Attribution project (2024). The fit allows for a dependence of the distribution parameters on explanatory variables, typically the GMST (e.g., Ciavarella et al., 2021; van Oldenborgh et al., 2021; Harrington et al., 2022; Philip et al., 2020). A GMST dependence is then interpreted as a response of extreme events to ACC. This dependence is then used to attribute extreme events by quantifying the effects of anthropogenic forcing on the probability of their occurrence. Technically, this is achieved by writing the location parameter of the extreme event distribution, $\mu$, as a function of $\overline{T}$, the (low-pass filtered, via a 4-year running average) GMST “as a measure of anthropogenic climate change” Harrington et al. (2022), such that $\mu=\mu_{0}+\alpha\overline{T}$. The GMST-dependence, $\alpha$, is estimated via a fit to observations. The fitted PDF is then used to evaluate the probability of extreme events at 1900, $P(\mu_{0}+\alpha\overline{T}(1900))$ relative to those at 2020, $P(\mu_{0}+\alpha\overline{T}(2020))$. This allows for estimating how the probability of an extreme event changed between these times. A change in the probability of an extreme event is inferred by the WWA approach if $\alpha$ is non-zero, based on the change in GMST.
Although WWA does not carry out formal hypothesis testing, we interpret that they adopt a null hypothesis of $\alpha$ equal to zero implying no ACC influence on extremes and an alternative hypothesis that, if $\alpha$ is not equal to zero, ACC does influence extremes. This issue is discussed in detail in section 22.6 and further addressed throughout the text.
Our focus is on the empirically-based approaches used by WWA, but it should be noted that WWA also requires evidence from climate model simulations before making any attribution statements. WWA often uses multiple climate models and is careful to exclude models based on their performance and return periods relative to observations (e.g., Pinto et al., 2023). In this study we use a single model (CESM) to demonstrate some of our points but focus on assessment of observational methodologies associated with fitting of extreme value distributions to observations. Our goal is limited to evaluating the observational approach described by WWA and suggesting additional tests to examine its robustness, as opposed to developing alternative attribution approaches.
@@
The issue highlighted by our analysis is simple: internal variability modes such as ENSO affect both local temperature extremes and GMST. In some cases, the effect on local extremes (say the effects of ENSO on heatwaves in Australia) may be large, but the effect on GMST is small. This might lead to a false conclusion that a small GMST signal leads to a comparatively large effect on local extremes. To demonstrate this point, we analyze a preindustrial climate model run with no anthropogenic forcing. We find, for example, that the fit of Australian heat extremes leads to values of $\alpha$ that are substantially greater than zero in the absence of ACC because of the effects of internal climate variability modes such as ENSO on both GMST and on local extremes. Our findings show that $\alpha$ can be strongly affected by internal variability, biasing the attribution results and complicating the WWA approach.
@@ 2.1 The Generalized Extreme Value (GEV) distribution
Extreme value distributions are a family of PDFs that can be used to represent the statistical behavior of block maxima or minima of a record. An example is an annual time series of the maximum daily temperature at each year, the “block” being one year in this case. It can be shown that the distribution of such maxima follows one of three classes of PDFs: Gumbel, Fréchet, or Weibull Coles et al. (2001), depending on the shape of the tail of the distribution of the events for which block maxima are calculated. Rather than explicitly specifying one of these three classes in fitting data for extreme events, one can combine these distributions into one functional form, represented by the Generalized Extreme Value (GEV) distribution. Large quantities of data are typically needed to accurately fit a GEV and calculate its parameters (e.g., Philip et al., 2020; Trevino et al., 2020).
@@ 2.2 Generalized Pareto Distribution (GPD)
The exponential fit was assumed by HW2022 in order to ensure that the scaling factor $\exp(\alpha\overline{T}/u_{0})$ applied to both $u_{0}$ and $\sigma_{0}$ is strictly positive and maintain a fixed ratio between the mean and variance. We obtain the optimized GPD parameters ($u_{0},\sigma,\xi,\alpha$) by again maximizing the log-likelihood (Eq. 4.10, Coles et al., 2001),
@@ 2.3 Uncertainty estimates and empirical cumulative distribution functions
We follow WWA studies and use non-parametric bootstrapping to estimate uncertainty ranges for the optimized parameters. In particular, we use the uncertainty in $\alpha$ as one way to evaluate the validity of the attribution results. We also compare the appropriately fitted distributions (GEV and normal distributions for extreme temperatures or GPD and log-normal for droughts) to the empirical cumulative distribution functions (CDFs) of the observations to assess the quality of the fit. The empirical CDF is calculated by sorting extreme values (either temperature or precipitation) from smallest to largest, calculating the cumulative sum, and normalizing to a maximum CDF value of one. We calculate error bars in the empirical CDFs based on bootstrap resampling with replacement of the observations. The empirical CDF is calculated for each sampling, and we then calculate the 5–95th and 2.5–97.5th percentile ranges from the resulting distribution at each value of the variable whose CDF is estimated. In order to estimate uncertainty in the parameters estimated from our GEV or GPD maximum likelihood fit, we similarly use bootstrap resampling. We sample with replacement the GMST and corresponding extreme temperature record 5,000 times and optimize the parameters for each resampling. We then estimate the 90th and 95th percentiles from 5,000 bootstrap samples. For the Madagascar precipitation case, we perform 5,000 bootstrap resamples when analyzing the observations, but 1,000 bootstrap resamples when analyzing the CESM Large Ensemble due to the slow convergence and expensive computations using the GPD log-likelihood maximization.
@@ 2.6 Hypothesis testing
Although WWA does not carry out formal hypothesis testing, we interpret that they adopt a null hypothesis of $\alpha$ equal to zero implying no ACC influence on extremes and an alternative hypothesis that, if $\alpha$ is not equal to zero, ACC does influence extremes. This inference is based on the quote, “If the one- or two-sided 95% confidence interval [for the change in probability, which has the same effect as $\alpha$] excludes zero, the change is statistically significant” (see Section 4.3.1 of Philip et al., 2020). We assess whether rejection of a null hypothesis of $\alpha$ equal to zero permits accepting an alternate of ACC influencing extreme events. In other words, we assess if a nonzero $\alpha$ necessarily means a role for ACC. We note that another null hypothesis for purposes of heat could be that the location of PDF of heat extremes is shifted by an amount equaling the GMST warming, as would be represented by $\alpha=1$. For purposes of consistency with the null implied by WWA, however, we only consider a null of $\alpha$ equal to zero.
WWA uses two variants of their analysis approach, one in which the extreme event in question is incorporated into the estimate of the PDFs, which we follow here, and one in which it is excluded Philip et al. (2020). The act of choosing to study a class of events immediately following the occurrence of a particularly extreme example is known to induce a bias related to whether it is included or excluded Miralles and Davison (2023). We choose to include the extreme value in question, which tends to make the parametric dependence on GMST larger and tends to make it easier to reject the null of no change in favor of the alternate hypothesis. Within the context of the WWA hypothesis test, this choice to include the extreme in question makes it more likely to conclude that AAC significantly influences extreme events.
@@ 2.7 Data
We also study GMST anomalies and extreme temperatures over the Siberian region defined in CC2021 within the CESM Large Ensemble Kay et al. (2015) over the period 1926–2019. The CESM Large Ensemble contains 40 simulations of the climate over the period 1920–2100 under historical and RCP 8.5 external forcing, of which we use the 35 that were run on the NCAR supercomputer. These simulations differ only in small perturbations to the initial conditions, which lead to different temporal sequences of modes of internal variability. We use model output only up to the year 2020, and, as a result, there is little sensitivity to the emissions scenario used (DAMIP, Gillett et al., 2016, uses RCP 4.5) because emissions scenarios only significantly diverge later. We do not use some additional runs done on the University of Toronto supercomputer that contain a global mean temperature bias National Center for Atmospheric Research (2024).
@@ 3 Results
We first describe our analyses of how extreme events depend on GMST in the context of the long CESM pre-industrial simulation in subsection 33.1. The simulation contains no time-varying anthropogenic forcing, making it useful for evaluating a null-hypotheses associated with no anthropogenic influence. We then consider each of the three case studies in subsections 33.2, 33.3, and 33.4. For each case we assess how GEV or GPD distributions fit the observations relative to normal or log-normal distributions. We then evaluate the justification for the inclusion of a GMST dependence of the distribution parameters using the deviance statistic and estimate the uncertainty level of the parameter $\alpha$ used for the attribution. Finally, we examine the amount of data needed to adequately constrain GMST dependence in a GEV or GPD fit.
@@ 3.1 Extremes analysis of a preindustrial simulation
WWA recognizes that $\alpha$ may be influenced by internal variability and, as noted, use a 4-year running mean GMST to filter out the effects of short-term internal variability. This approach, however, may not fully remove the ENSO signal. Furthermore, there is also the potential for nonlinear interactions between ENSO and GMST as well as the presence of slower modes of variability such as the Pacific Decadal Oscillation (PDO).
@@
A similar analysis of Australian heat extremes made using the CESM Large Ensemble, which includes anthropogenic forcing, gives a distribution of $\alpha$ that is closer to zero (blue bars in Fig. 1) than in the unforced simulations (green bars). This indicates that both ACC and internal variability influence regional extremes and GMST, such that estimates of $\alpha$ will reflect multiple processes. In simulations with time-varying anthropogenic forcing, however, there is no way to disentangle the effect of internal variability on $\alpha$ using the WWA approach as described in published literature. We conclude that the magnitude of $\alpha$ may be a biased indicator of anthropogenic influence. Including indices of internal variability as covariates (by adding parameters that represent the dependence of the distribution on regional climate variability modes such as ENSO) may help to disentangle their respective contributions. Yet the projection of internal variability onto GMST can still confound the results, and adding additional parameters to the distribution to represent the dependence of the distribution on regional climate variability modes such as ENSO may not be justified given that below we find that even adding the single parameter $\alpha$ can be questioned due to the short observed record.
@@
These results highlight a potential major difficulty with analyses based on a fit of distribution parameters to GMST in a more realistic scenario that includes both internal variability and ACC: the calculated $\alpha$, and therefore the shift in the probability of extreme events, may be due to internal variability rather than only ACC. When analyzing the observational record in section 3.3, we find $\alpha\approx 2\;^{\circ}$ per C. This suggests a two-degree warming of extreme temperatures for a one-degree warming of the GMST, which is at the upper end of the range of values estimated from CESM LE. Based on our results from the CESM preindustrial run with a time-invariant anthropogenic forcing, we therefore suspect that the estimate obtained for $\alpha$ from the observed record is biased high by internal variability.
@@
Analysis of the CESM preindustrial run for the case of Siberian heatwaves (green bars in Fig. 1b) shows that the value for $\alpha$ deduced from a record similar in length to that of the observations ranges between 1 to 4, although there is again no time-varying anthropogenic forcing in the corresponding model run. This large and variable response implies that a non-zero $\alpha$ does not necessarily imply an effect of ACC. It may suggest that natural variability affects the value of $\alpha$ for Siberian heat waves during some decades but not others. This analysis shows that $\alpha$ equal to zero is an unsuitable null hypothesis for the purposes of evaluating the presence of an anthropogenic influence on extreme events (for context, we note again that WWA does not explicitly define a null hypothesis).
@@
The result that internal variability may strongly affect $\alpha$ when the 4-year average GMST is used to link extreme events to ACC suggests that it might be useful to reevaluate some of the WWA results based on the fit to extreme distributions. It is possible, of course, that a strong anthropogenic signal in a few decades could significantly affect extreme events and overwhelm the effects of internal variability in cases such as the Siberian Heatwave. The attribution of extreme temperatures may be more difficult in cases like the Australian example, where the effects of internal variability are large, as indicated by the large $\alpha$ detected in the CESM preindustrial simulation. In such cases, it would be important to correctly identify all relevant covariates — such as the possible effect of ENSO in the Australian case — and note that misleading results for $\alpha$ may be obtained if one relies only on GMST.